I need help solving 3c+2d=2
D=4
By substitution

Answer by algebrahouse.com(1659)

3c + 2d = 2
d = 4

3c + 2(4) = 2 {substituted 4 for d}
3c + 8 = 2 {multiplied}
3c = -6 {subtracted 8 from each side}
c = -2 {divided each side by 3}

c = -2 and d = 4

For more help from me, visit: www.algebrahouse.com